OpenStudy (anonymous):

A right triangle has base (x - 7) units and height (2x - 10) units. Part A: What is the square of the length of the hypotenuse of the triangle? Show your work. (4 points) Part B: What is the area of the triangle? (3 points) Part C: Using the solution obtained in Part B, explain the closure property of multiplication of polynomials. (3 points)

OpenStudy (anonymous):

@ganeshie8 @Nnesha @hartnn

OpenStudy (anonymous):

Please help Anyone

OpenStudy (anonymous):

@mathmate @welshfella Can you guys help? I have a general idea of what I need to do for part one, but I get really lost afterwards.

OpenStudy (welshfella):

square of hypotenuse , by Pythagoras theorem = sum of squares on the other 2 sides = (x - 7^2 + (2x - 10)^2

OpenStudy (anonymous):

I got that much, but I need to simplify it I believe

OpenStudy (anonymous):

@welshfella

OpenStudy (kobeni-chan):

What do you have that you need to simplify?

OpenStudy (anonymous):

I need the square length of the hypotenuse. so wouldn't I need to simplify (x-7)^2 + (2x+10)= C^2?

OpenStudy (anonymous):

Like I know that much but I don't know what to do afterwards. do I square everything?

OpenStudy (kobeni-chan):

You need to square (x-7) and (2x-10) right now. If you have a scientific calculator you can do this easily, but if you have a reg. calculator then do (x-7)(x-7) and use the distributive property. Then do the same with (2x-10)(2x-10)

OpenStudy (anonymous):



OpenStudy (anonymous):